### v1

# New rigidity theorem of Einstein manifolds and curvature operator of the second kind

Abstract.
Using Bochner techniques, we prove that a compact Einstein manifold of dimension $n\geq 4$ has constant curvature provided that the curvature operator of the second kind satisfies a cone condition that is strictly weaker than nonnegativity. Furthermore, employing a result of Li [Li5], we establish that any closed Einstein manifold of dimension $n\geq 4$ satisfying
$k^{-1}({\lambda}_{1}+\cdots+{\lambda}_{k})\geq-\theta(n,k)\bar{\lambda},\quad\text{for some}\quad k\leq[\frac{n+2}{4}]$
must be either flat or a spherical space form. Here, ${\lambda}_{1}\leq{\lambda}_{2}\leq\cdots\leq{\lambda}_{\frac{(n-1)(n+2)}{2}}$ are the eigenvalues of $\mathring{R}\,$, $\bar{\lambda}$ is their average, and $\theta(n,k)$ is a positive constant. This result generalizes the work of Dai-Fu [DF] and Chen-Wang [CW1, CW].
Key words and phrases:
Einstein manifolds, Bochner technique, Sphere theorems
2020 Mathematics Subject Classification:
53C20, 53C24, 53C25.
Supported in part by National Natural Science Foundations of China #12461008 and 12271069, Jiangxi Province Natural Science Foundation of China #20202ACB201001, Jiangxi Province Graduate Student Innovation Special Fund Project #YC2025-B037.
1. introduction
The study of the Riemannian curvature operator of the second kind originated with Nishikawa’s conjecture [Ni], which states that a closed simply connected Riemannian manifold with a nonnegative curvature operator of the second kind is diffeomorphic to a locally symmetric space, and if the operator is positive, the manifold is diffeomorphic to a spherical space form. This conjecture was confirmed by Cao–Gursky–Tran [CGT] under the assumption of two-positivity, and later strengthened by Li [Li4] and Nienhaus–Petersen–Wink [NP] under the weaker condition of three-nonnegativity. Further classification results under various curvature conditions of the second kind can be found in [DFD, Li3, Li7, Li6, Li4, Li2, Li5, DF, NP].
[…]
Recently, Li [Li5] studied cone conditions of the curvature operator of the second kind, which satisfies
${\alpha}^{-1}({\lambda}_{1}+\cdots+{\lambda}_{\alpha})\geq-\theta\bar{\lambda},$
for its eigenvalue of curvature operator of the second kind. This condition is denote by $C(\alpha,\theta)$ and $\mathring{C}\,(\alpha,\theta)$ if the inequality is strict. Li proved that if $\mathring{R}\,\in C(\frac{n+2}{2},\theta)$ for some $1<\theta<\frac{2(n-1)}{n+2},$ then $M$ is either flat or a rational homology sphere.
Building on the work of Li [Li5] and Dai–Fu [DF], Cheng and Wang [CW] relaxed the curvature assumption to a cone condition. They introduced a positive constant $\theta(n)$ and proved that if an $n$-dimensional closed Einstein manifold ($n=4,5$, or $n\geq 8$) satisfies $\lambda_{1}+\lambda_{2}\geq-2\theta(n)\bar{\lambda}$, then it is either flat or a spherical space form. This condition is strictly weaker than $2$-nonnegativity, which corresponds to $\theta=0$.
In this paper, we investigate Einstein manifolds of dimension $n\geq 4$ whose curvature operator of the second kind satisfies a cone condition that is strictly weaker than nonnegativity. Let $\{\lambda_{i}\}_{i=1}^{N}$ be the eigenvalues of $\mathring{R}\,$ and $\bar{\lambda}=\sum\nolimits_{i=1}^{N}{{\lambda}_{i}}/{N}\;$ their average. By applying results of Dai–Fu [DF] and Li [Li5], we obtain the following theorems.
Theorem A.
Let $(M^{n},g)$ be an $n$-dimensional compact Einstein manifold. For $n\geq 8,n\neq 10$ and $k\leq[\frac{n+2}{4}]$, there exists a constant $\theta(n,k)>0$ such that if the curvature operator of the second kind satisfies
$k^{-1}({\lambda}_{1}+\cdots+{\lambda}_{k})\geq-\theta(n,k)\bar{\lambda},$
where
$\theta(n,k)=\frac{(N-k)(2N-9n+6)-3nN(k-2)}{(N-3)(N-k)+3kn(N-2)},N=\frac{(n-1)(n+2)}{2},$
then $M$ is either flat or a spherical space form.
When $n=4,5$ we have the following theorem.
Theorem B.
Let $(M,g)$ be an $n=4$ or $5$-dimensional compact Einstein manifold, and $\mathring{R}\,$ be the curvature operator of the second kind. If
(i) $n=4$ and $\mathring{R}\,$ satisfies ${\lambda}_{1}\geq-\frac{17}{7}\bar{\lambda}$ or ${\lambda}_{1}+{\lambda}_{2}\geq-\bar{\lambda}$, or
(ii) $n=5$ and $\mathring{R}\,$ satisfies ${\lambda}_{1}\geq-\frac{47}{18}\bar{\lambda}$ or ${\lambda}_{1}+{\lambda}_{2}\geq-\frac{4}{3}\bar{\lambda}$ or ${\lambda}_{1}+{\lambda}_{2}+{\lambda}_{3}\geq-\frac{1}{18}\bar{\lambda}$,
then $M$ is either flat or a spherical space form.
Remark 1.1.
For dimensions $n=6,7$, we have $\frac{n+2}{4}]=2$, yet our method does not yield a corresponding cone condition for $k=2$. Similarly, for $n=10$, the method does not provide a cone condition for $k=3$. Nevertheless, the following theorem remains valid for the cases $n=6,7$ with $k=1$, and for $n=10$ with $k=1,2$.
Theorem C.
Let $(M,g)$ be an $n=6,7$ or $10$-dimensional compact Einstein manifold, and $\mathring{R}$ be the curvature operator of the second kind. If
(i) $n=6$ and $\mathring{R}\,$ satisfies ${\lambda}_{1}\geq-\frac{208}{647}\bar{\lambda}$,
(ii) $n=7$ and $\mathring{R}\,$ satisfies ${\lambda}_{1}\geq-\frac{163}{383}\bar{\lambda}$, or
(iii) $n=10$ and $\mathring{R}\,$ satisfies ${\lambda}_{1}\geq-\frac{964}{1421}\bar{\lambda}$ or ${\lambda}_{1}+{\lambda}_{2}\geq-\frac{16}{37}\bar{\lambda},$
then $M$ is either flat or a spherical space form.
Remark 1.2.
When $\theta=0$, Theorem A,B,C are the result of Dai-Fu [DF]. When $k=1,2$, Theorem A,B,C are the result of Chen-Wang [CW1, CW].
2. notation and lemmas
Here, we introduce our notation and summarize key facts regarding the curvature operator of the second kind. Further details can be found in [Li4, Li5, NP, DF, CW].
Let $(V,g)$ be an $n$-dimensional Euclidean vector space and $\left\{e_{i}\right\}_{i=1}^{n}$ be an orthonormal basis for $V$. We recall that the second-order tensor space can be decomposed by
$T^{2}(V)={\Lambda}^{2}(V)\oplus S^{2}(V),$
where ${\Lambda}^{2}(V)$ and $S^{2}(V)$ is the second-order anti-symmetric tensor space and the second order symmetric tensor space on $V$, respectively. The space $S^{2}(V)$ can split subspaces as
$S^{2}(V)=S_{0}^{2}(V)\oplus\mathbb{R}g,$
where $S_{0}^{2}(V)$is the space of traceless symmetric two-tensors, and $g=\sum\limits_{i=1}^{n}{e_{i}\otimes e_{i}}$. We always denote by $N=\frac{(n-1)(n+2)}{2}$ the dimension of $S_{0}^{2}(V)$ in this paper.
The space of symmetric two-tensors on ${{\Lambda}^{2}}(V)$ has the orthogonal decomposition
$S^{2}({\Lambda}^{2}(V))=S_{B}^{2}({\Lambda}^{2}(V))\oplus{\Lambda}^{4}(V),$
where $S_{B}^{2}({\Lambda}^{2}(V))$ consists of all tensors $R\in S^{2}({\Lambda}^{2}(V))$ that also satisfy the first Bianchi identity. $S_{B}^{2}({\Lambda}^{2}(V))$ is called the space of algebraic curvature operators. By the symmetries of $R\in S_{B}^{2}({{\Lambda}^{2}}(V))$, there are (up to sign) two natural ways in which induces a symmetric linear map from $T^{2}(V)$ to $T^{2}(V)$. The first one is defined by
$\hat{R}:{{\Lambda}^{2}}(V)\to{{\Lambda}^{2}}(V),$
$\hat{R}({{e}_{i}}\wedge{{e}_{j}})=\frac{1}{2}\sum\limits_{k,l}{{{R}_{ijkl}}{{e}_{k}}\wedge{{e}_{l}}},$
which is called the curvature operator of the first kind. The second one denoted by $\bar{R}$, is defined as
$\bar{R}:{{S}^{2}}(V)\to{{S}^{2}}(V),$
$\bar{R}({{e}_{i}}\odot{{e}_{j}})=\sum\limits_{k,l}{{{R}_{iklj}}{{e}_{k}}\odot{{e}_{l}}},$
$\bar{R}$ induces a symmetric form $\mathring{R}\,$ defined by
$\mathring{R}\,:S_{0}^{2}(V)\times S_{0}^{2}(V)\to R$
$({{S}^{\alpha}},{{S}^{\beta}})\to{{R}_{ijkl}}S_{jk}^{\alpha}S_{il}^{\beta},$
which is called curvature operator of the second kind. It was introduced by Nishikawa [Ni] and as pointed out in [NP], the curvature operator of the second kind $\mathring{R}\,$ can also be interpreted as $\bar{R}$ and its image restrict on $S_{0}^{2}(V)$.
Let $(M,g)$ be an $n$-dimensional Riemannian manifold. For each $p\in M,$ let $V=T_{p}M$ and $R$ is denoted by the Riemannian curvature tensor of $M$. $\mathring{R}\,$ is denoted by curvature operator of second kind of $M$. We introduce the following definition which comes from [NP]
Definition 2.1.
Let $T^{(0,k)}(V)$denote the space of $(0,k)$-tensor space on $V$. For $S\in S^{2}(V)$ and $T\in T^{(0,k)}(V),$ we define
$S:T^{(0,k)}(V)\to T^{(0,k)}(V)$
$(ST)(X_{1},\cdots,X_{k})=\sum\limits_{i=1}^{k}{T(X_{1},\cdots,SX_{i},\cdots,X_{k})}$
and define $T^{S^{2}}\in T^{(0,k)}(V)\otimes S^{2}(V)$ by
$\left\langle T^{S^{2}}(X_{1},\cdots,X_{k}),S\right\rangle=(ST)(X_{1},\cdots,X_{k}).$
Hence, if $\left\{{\tilde{S}}^{i}\right\}_{i=1}^{\frac{n(n+1)}{2}}$ is an orthonormal basis for $S^{2}(V)$, then
$T^{S^{2}}=\sum\limits_{i=1}^{\frac{n(n+1)}{2}}{{\tilde{S}}^{i}T\otimes{\tilde{S}}^{i}}.$
Similarly, we define $T^{S_{0}^{2}}\in T^{(0,k)}(V)\otimes S_{0}^{2}(V)$ and for an orthonormal basis $\left\{S^{i}\right\}$ of $S_{0}^{2}(V)$
$T^{S_{0}^{2}}=\sum\limits_{i}{S^{i}T\otimes S^{i}}.$
Let $\{{\lambda}_{i}\}$ denote the eigenvalues of the operator $\mathring{R}\,$ and let $\left\{S^{i}\right\}$ be the corresponding orthonormal eigenbasis of $S_{0}^{2}(V)$. In what follows, the eigenvalues $\{{\lambda}_{i}\}$ are assumed to be arranged in non‑decreasing order. Define the action of $\mathring{R}\,$ on the space $T^{S_{0}^{2}}$ as
$\mathring{R}\,(T^{S_{0}^{2}})=\sum\limits_{i}{S^{i}T\otimes\mathring{R}\,(S^{i}),}$
then
$\left\langle\mathring{R}\,(T^{S_{0}^{2}}),T^{S_{0}^{2}}\right\rangle=\sum\limits_{i}{{\lambda}_{i}\left|S^{i}T\right|^{2}}.$
Let $W$ be the Weyl curvature tensor of $M$, Set $S=\sum\limits_{i=1}^{N}{{{\left|{{S}^{i}}W\right|}^{2}}}$, $M=\underset{1\leq i\leq N}{\mathop{\max}}\,{{\left|{{S}^{i}}W\right|}^{2}}.$ The following estimate holds for the cone condition.
Proposition 2.2.
Let ${\lambda}_{1}\leq{\lambda}_{2}\leq\cdots\leq{\lambda}_{N}$ with $\sum\limits_{i=1}^{N}{{\lambda}_{i}}=N\bar{\lambda}.$ Under the cone condition $k^{-1}({\lambda}_{1}+\cdots+{\lambda}_{k})\geq-\theta\bar{\lambda}$ for $0<k\leq[\frac{S}{M}]$, we have
$\sum\limits_{i=1}^{N}{{\lambda}_{i}\left|S^{i}W\right|^{2}}\geq-S\theta\bar{\lambda}.$
Proof.
First we have
$\sum\limits_{i=1}^{N}{{\lambda}_{i}\left|S^{i}W\right|^{2}}$ $\geq$ $\sum\limits_{i=1}^{k}{{\lambda}_{i}\left|S^{i}W\right|^{2}}+\sum\limits_{i=k+1}^{N}{{\lambda}_{k+1}\left|S^{i}W\right|^{2}}$ (2.1)
$=$ ${{\lambda}_{k+1}}\sum\limits_{i=1}^{N}{\left|S^{i}W\right|^{2}}+\sum\limits_{i=1}^{k}{({\lambda}_{i}-{\lambda}_{k+1})\left|S^{i}W\right|^{2}}$
$\geq$ ${\lambda}_{k+1}S+M\sum\limits_{i=1}^{k}{({\lambda}_{i}-{\lambda}_{k+1})}$
$=$ $(S-kM){\lambda}_{k+1}+M\sum\limits_{i=1}^{k}{\lambda}_{i},$
If $k^{-1}({\lambda}_{1}+\cdots+{\lambda}_{k})\geq-\theta\bar{\lambda}$, then
${{\lambda}_{k+1}}\geq\frac{{\lambda}_{1}+\cdots+{\lambda}_{k}}{k}\geq-\theta\bar{\lambda}.$ (2.2)
Applying inequality (2.2) to the right-hand side of (2.1), we get for $0<k\leq[\frac{S}{M}],$
$\sum\limits_{i=1}^{N}{{\lambda}_{i}\left|S^{i}W\right|^{2}}$ $\geq$ $(S-kM){\lambda}_{k+1}+M\sum\limits_{i=1}^{k}{\lambda}_{i}$
$\geq$ $-\theta\bar{\lambda}(S-kM)-Mk\theta\bar{\lambda}$
$=$ $-S\theta\bar{\lambda}.$
3. proof of main theorem
Let $(M^{n},g)$ be an Einstein manifold of dimension $n$. Denote by $\{\lambda_{i}\}_{i=1}^{N}$ the eigenvalues of the operator $\mathring{R}\,$ and let $\bar{\lambda}=\sum\nolimits_{i=1}^{N}{{\lambda}_{i}}/{N}\;$ be their average. From [Li4], it follows that
$tr(\mathring{R})=\frac{n+2}{2n}s,$
where $s$ is the scalar curvature of $M$. Consequently $\bar{\lambda}=\frac{s}{n(n-1)}$.
Based on the calculations presented in [DF], it has been calculated that
$S=\sum\limits_{i=1}^{N}{{{\left|{{S}^{i}}W\right|}^{2}}}=\frac{2({{n}^{2}}+n-8)}{n}{{\left|W\right|}^{2}}=\frac{4N-12}{n}{{\left|W\right|}^{2}},$ (3.1)
and
$M=\underset{1\leq i\leq N}{\mathop{\max}}\,{\left|S^{i}W\right|^{2}}=\frac{8n-16}{n}{{\left|W\right|}^{2}}.$
Hence
$\frac{S}{M}]\geq[\frac{n+2}{4}].$
Moreover, on Einstein manifold, the square norm of the Weyl curvature can be expressed as
${{\left|W\right|}^{2}}=\frac{4}{3}\sum\limits_{i=1}^{N}{\lambda_{i}^{2}}-\frac{4N}{3}{{\bar{\lambda}}^{2}}.$ (3.2)
With these preparations we obtain the following estimate for the Bochner formula.
Lemma 3.1.
Let $(M^{n},g)$ be an Einstein manifold of dimension $n\geq 6$. If $k^{-1}({\lambda}_{1}+\cdots+{\lambda}_{k})\geq-\theta\bar{\lambda}$ for $k\leq[\frac{n+2}{4}],\theta\geq 0$, then the curvature operator $R$ of $M$ satisfies
$3\left\langle\Delta R,R\right\rangle$ $\geq$ $16\sum\limits_{i=1}^{N}{\lambda_{i}^{3}}+\frac{16}{3n}[2N-12n+6-(N-3)\theta]\bar{\lambda}\sum\limits_{i=1}^{N}{\lambda_{i}^{2}}$ (3.3)
$-\frac{16N}{3n}[2N-9n+6-(N-3)\theta]\bar{\lambda}^{3}.$
Proof.
The following Bochner formula comes from Proposition 3.4 in [DF].
$3\left\langle\Delta R,R\right\rangle=$ $\sum\limits_{i}{{\lambda}_{i}\left|S^{i}W\right|^{2}}+8\left(\frac{-n^{3}+6n^{2}+12n-8}{3n^{4}(n-1)^{2}}\right)s^{3}$
$+8\left(\frac{2n^{2}-22n+8}{3n^{2}(n-1)}\right)s\sum\limits_{i}{\lambda_{i}^{2}}+16\sum\limits_{i}{\lambda_{i}^{3}}.$
Substituting $N=\frac{(n-1)(n+2)}{2}$ and $\bar{\lambda}=\frac{s}{n(n-1)}$, we simplify the equation to
$3\left\langle\Delta R,R\right\rangle$ $=$ $\sum\limits_{i=1}^{N}{{{\lambda}_{i}}{{\left|{{S}^{i}}W\right|}^{2}}}-\frac{16N(2N-9n+6)}{3n}{{{\bar{\lambda}}}^{3}}$
$+\frac{16(2N-12n+6)}{3n}\bar{\lambda}\sum\limits_{i=1}^{N}{\lambda_{i}^{2}}+16\sum\limits_{i=1}^{N}{\lambda_{i}^{3}}.$
Applying Proposition 2.2, together with (3.1) and (3.2) yields
$\sum\limits_{i=1}^{N}{{\lambda}_{i}\left|S^{i}W\right|^{2}}$ $\geq-\frac{4N-12}{n}\theta\bar{\lambda}\left|W\right|^{2}$
$=-\frac{16(N-3)}{3n}\theta\bar{\lambda}\sum\limits_{i=1}^{N}{{\lambda}_{i}^{2}}+\frac{16N(N-3)}{3n}\theta\bar{\lambda}^{3}.$
Combining the above two equations yields the desired result. ∎
Remark 3.2.
When $k=2$, this is Lemma 3.3 of [CW].
To establish the main theorem, we present the following algebraic results, which play a crucial role in the proof.
Lemma 3.3.
Let $0\leq{\lambda}_{1}\leq{\lambda}_{2}\leq\cdots\leq{\lambda}_{N}$ and $\sum\limits_{i=1}^{N}{{\lambda}_{i}}=C>0$. If $C=\frac{N(N-1)}{2N-1},$ then the function
$F(\lambda_{1},\lambda_{2},\ldots,\lambda_{N})=\sum\limits_{i=1}^{N}{\lambda_{i}^{3}}-\sum\limits_{i=1}^{N}{\lambda_{i}^{2}}$
attains its minimal at $\left(\frac{C}{N},\frac{C}{N},\ldots,\frac{C}{N}\right)$ and $\left(0,\frac{C}{N-1},\ldots,\frac{C}{N-1}\right).$
Proof.
We will divide the proof into two cases.
Case 1. Consider ${\lambda}_{1}>0$. Under the constraints $\sum\limits_{i}{{{\lambda}_{i}}}=C>0$, Lagrange multipliers yields a constant $\mu\in\mathbb{R}$ such that
$3\lambda_{i}^{2}-2{\lambda}_{i}+\mu=0$ (3.4)
for $i=1,\ldots,N.$ The possible solution of this equation is
$a=\frac{1-\sqrt{1-3\mu}}{3},b=\frac{1+\sqrt{1-3\mu}}{3}.$
Hence, the possible critical points are
$Q_{k}=\left(\overset{k}{\mathop{\overset{\frown}{a,\ldots,a}}}\,,\overset{N-k}{\mathop{\overset{\frown}{b,\ldots,b}}}\,\right).$
Under constrain ka+(Nk)b=Cka+(N-k)b=C, we obtain
$(N-2k)\sqrt{1-3\mu}=3C-N.$
Because $C=\frac{N(N-1)}{2N-1}$ implies 3CN>03C-N>0, we must have N2k>0N-2k>0. Consequently,
$\mu=\frac{1}{3}[1-{\left(\frac{3C-N}{N-2k}\right)^{2}}].$ (3.5)
Substituting $Q_{k}$ into $F$ and applying the relation (3.4) to obtain
$F(Q_{k})$ $=-\frac{C}{3}\mu-\frac{1}{3}\sum\limits_{i}{\lambda_{i}^{2}}$
$=-\frac{C}{3}\mu-\frac{N(2-3\mu)+2(N-2k)\sqrt{1-3\mu}}{27}.$
Together with (3.5) given
$F(Q_{k})$ $=-\frac{C}{3}\frac{1}{3}[1-{{\left(\frac{3C-N}{N-2k}\right)}^{2}}]-\frac{N}{27}-\frac{N{{\left(\frac{3C-N}{N-2k}\right)}^{2}}+2(3C-N)}{27}$
$=(\frac{C}{9}-\frac{N}{27}){{\left(\frac{3C-N}{N-2k}\right)}^{2}}-\frac{C}{3}+\frac{N}{27}.$
Since we assume $C=\frac{N(N-1)}{2N-1}>\frac{N}{3},$ it follow that $\frac{C}{9}-\frac{N}{27}>0$. Hence
$\min F(Q_{k})=\underset{0\leq k<{N}/{2}\;}{\mathop{\min}}\,\{(\frac{C}{9}-\frac{N}{27}){\left(\frac{3C-N}{N-2k}\right)^{2}}-\frac{C}{3}+\frac{N}{27}\}$
$=F(Q_{0}),$
and the minimal point is $\left(\frac{C}{N},\frac{C}{N},\ldots,\frac{C}{N}\right).$
Case 2, Consider ${\lambda}_{1}=0$. Suppose $X$ be a minimizer such that it has the maximum number of zeros among all minimizers. We claim that the non-zero elements of $X$ are all equal.
Assume, for a contradiction, that ${\lambda}_{i}>{\lambda}_{j}>0$ for some $i,j$. If ${\lambda}_{i}+{\lambda}_{j}\leq\frac{2}{3},$ we have
$\lambda_{i}^{3}-\lambda_{i}^{2}+\lambda_{j}^{3}-\lambda_{j}^{2}=({\lambda}_{i}+{\lambda}_{j})^{3}-({\lambda}_{i}+{\lambda}_{j})^{2}+{\lambda}_{i}{\lambda}_{j}(2-3{\lambda}_{i}-3{\lambda}_{j})$
$\geq({\lambda}_{i}+{\lambda}_{j})^{3}-({\lambda}_{i}+{\lambda}_{j})^{2}.$
Thus, replacing ${\lambda}_{i},{\lambda}_{j}$ with ${\lambda}_{i}+{\lambda}_{j},0$ yields another minimizer. This contradicts the fact that $X$ has the largest number of zeros.
If ${\lambda}_{i}+{\lambda}_{j}>\frac{2}{3},$ letting $u=\frac{{\lambda}_{i}+{\lambda}_{j}}{2}$, we have
$\lambda_{i}^{3}-\lambda_{i}^{2}+\lambda_{j}^{3}-\lambda_{j}^{2}=2(u^{3}-u^{2})+\frac{1}{4}(3{\lambda}_{i}+3{\lambda}_{j}-2)({\lambda}_{i}-{\lambda}_{j})^{2}$
$>2(u^{3}-u^{2}).$
Thus replacing ${\lambda}_{i},{\lambda}_{j}$ with $\frac{{\lambda}_{i}+{\lambda}_{j}}{2}$, $\frac{{\lambda}_{i}+{\lambda}_{j}}{2}$, yields a smaller minimizer, this contradicts the minimality of $X$. The claim is proved.
Through the above discussion, we can assume that for $X$, $0={\lambda}_{1}=\cdots={\lambda}_{k},$ and ${\lambda}_{k+1}=\cdots={\lambda}_{N}$ for some $1\leq k<N,$ then
$F(X)=\frac{C^{3}}{(N-k)^{2}}-\frac{C^{2}}{N-k}.$
So in $\partial\Omega,$
$\underset{\partial\Omega}{\mathop{\min}}\,F=\underset{1\leq k<N}{\mathop{\min}}\,\frac{C^{3}}{(N-k)^{2}}-\frac{C^{2}}{N-k}.$
Since we assume $C=\frac{N(N-1)}{2N-1},$ we obtain
$\frac{C^{3}}{(N-k)^{2}}-\frac{C^{2}}{N-k}-(\frac{C^{3}}{N^{2}}-\frac{C^{2}}{N})=\frac{{C^{2}}(N^{2}-(N-k)^{2})(C-\frac{N(N-k)}{2N-k})}{(N-k)^{2}N^{2}}\geq 0,$
equality holds if and only if $k=1$. Hence,
$\underset{\partial\Omega}{\mathop{\min}}\,F=\underset{1\leq k<N}{\mathop{\min}}\,\frac{C^{3}}{(N-k)^{2}}-\frac{C^{2}}{N-k}=\frac{C^{3}}{N^{2}}-\frac{C^{2}}{N},$
and the minimal point is $\left(0,\frac{C}{N-1},\ldots,\frac{C}{N-1}\right)$. As $X$ possesses at least one zero, it follows that the minimizer must be unique.
Combining case 1 and case 2, we have completed the proof of the Lemma. ∎
Remark 3.4.
From case 2, we can conclude that if $C>\frac{N(N-1)}{2N-1}$, then the only minimal point of $F$ is $\left(\frac{C}{N},\frac{C}{N},\ldots,\frac{C}{N}\right)$.
When $k^{-1}({\lambda}_{1}+{\lambda}_{2}+\cdots+{\lambda}_{k})\geq-\theta\bar{\lambda}$ for some constant $\theta$, the following estimate holds for ${\lambda}_{1}$.
Lemma 3.5.
Let ${\lambda}_{1}\leq{\lambda}_{2}\leq\cdots\leq{\lambda}_{N}$ be a sequence with ${\lambda}_{1}+{\lambda}_{2}+\cdots+{\lambda}_{N}=N\bar{\lambda}$. If there is constant $\theta$ such that $k^{-1}({\lambda}_{1}+{\lambda}_{2}+\cdots+{\lambda}_{k})\geq-\theta\bar{\lambda}$ for some $k<N$, then
${{\lambda}_{1}}\geq-\frac{(N-1)k\theta+N(k-1)}{N-k}\bar{\lambda}.$
Proof.
Since ${\lambda}_{i}$} is a non-decreasing sequence, and $k^{-1}({\lambda}_{1}+{\lambda}_{2}+\cdots+{\lambda}_{k})\geq-\theta\bar{\lambda}$,
${{\lambda}_{i_{1}}}+{{\lambda}_{i_{2}}}+\cdots+{{\lambda}_{i_{k}}}\geq-k\theta\bar{\lambda}\quad\text{for any}\quad 1\leq i_{1}<i_{2}<\cdots<i_{k}\leq N.$
Among all such summation sequences, there are $\left(\begin{aligned} &N-1\\ &k-1\\ \end{aligned}\right)$ of them that contain ${\lambda}_{1}$, summing over these sequences obtain
$\left(\begin{aligned} &N-1\\ &k-1\\ \end{aligned}\right){{\lambda}_{1}}+\left(\begin{aligned} &N-2\\ &k-2\\ \end{aligned}\right)\sum\limits_{i=2}^{N}{{\lambda}_{i}}\geq-\left(\begin{aligned} &N-1\\ &k-1\\ \end{aligned}\right)k\theta\bar{\lambda}$
since there are $\left(\begin{aligned} &N-2\\ &k-2\\ \end{aligned}\right)$ sequences that include ${\lambda}_{1}$ and ${\lambda}_{i},i\neq 1.$ For ${{\lambda}_{1}}+{{\lambda}_{2}}+\cdots+{{\lambda}_{N}}=N\bar{\lambda}$, we finally get
$\left[\left(\begin{aligned} &N-1\\ &k-1\\ \end{aligned}\right)-\left(\begin{aligned} &N-2\\ &k-2\\ \end{aligned}\right)\right]{{\lambda}_{1}}\geq-\left[\left(\begin{aligned} &N-1\\ &k-1\\ \end{aligned}\right)k\theta+N\left(\begin{aligned} &N-2\\ &k-2\\ \end{aligned}\right)\right]\bar{\lambda},$
i.e. ${{\lambda}_{1}}\geq-\frac{(N-1)k\theta+N(k-1)}{N-k}\bar{\lambda}.$
Proof of Theorem A.
If $\bar{\lambda}=0$, then $M$ is flat. Now we consider $\bar{\lambda}>0$.
Writing the right term of (3.3) as a function $F(k,\theta,\lambda)$,
$F(k,\theta,\lambda)=16\sum\limits_{i=1}^{N}{\lambda_{i}^{3}}+\frac{16}{3n}(2N-12n+6-(N-3)\theta)\bar{\lambda}\sum\limits_{i=1}^{N}{\lambda_{i}^{2}}-\frac{16N}{3n}(2N-9n+6-(N-3)\theta)\bar{\lambda}^{3}.$
Now we consider the minimal of $F$ under constraint
${\lambda}_{1}\leq{\lambda}_{2}\leq\cdots\leq{\lambda}_{N},\quad k^{-1}({\lambda}_{1}+\cdots+{\lambda}_{k})\geq-\theta\bar{\lambda}\quad\text{and}\quad\sum\limits_{i=1}^{N}{{\lambda}_{i}}=N\bar{\lambda}.$
Obviously, $F(k,\theta,\bar{\lambda})=0$ when ${\lambda}_{1}={\lambda}_{2}=\cdots{\lambda}_{N}=\bar{\lambda}$, since in this case $M$ has constant curvature and the estimate in Lemma 3.1 becomes trivial.
Set
${\beta}_{i}={\lambda}_{i}+D\bar{\lambda},\quad\text{where}\quad D=\frac{(N-1)k\theta+N(k-1)}{N-k}.$
From Lemma 3.5 we know
${\beta}_{i}\geq 0\quad\text{and}\quad\sum\limits_{i}{{\beta}_{i}}=N(1+D)\bar{\lambda},$
then $F(k,\theta,\lambda)$ will become
$F(k,\theta,\beta)=$ $16\sum\limits_{i=1}^{N}{({\beta}_{i}-D\bar{\lambda})^{3}}+\frac{16}{3n}(2N-12n+6-(N-3)\theta)\bar{\lambda}\sum\limits_{i=1}^{N}{({\beta}_{i}-D\bar{\lambda})^{2}}$
$-\frac{16N}{3n}(2N-9n+6-(N-3)\theta){\bar{\lambda}}^{3}$
$=$ $16\left\{\sum\limits_{i=1}^{N}{\beta_{i}^{3}}+[\frac{1}{3n}(2N-12n+6-(N-3)\theta)-3D]\bar{\lambda}\sum\limits_{i=1}^{N}{\beta_{i}^{2}}\right\}+G(\theta,k){\bar{\lambda}}^{3}.$
where $G(\theta,k)$ is a polynomial in terms of $\theta$ and $k$ and it is independent of the variables ${\beta}_{i}$. When $F$ is regarded as a function of the the term ${\beta}_{i}$, $G(\theta,k)$ is therefore a constant and can be omitted specific expressions, as it does not influence the minimizer of $F$.
We choose $\theta$ such that
$\frac{1}{3n}(2N-12n+6-(N-3)\theta)-3D=-\frac{2N-1}{N-1}(1+D).$
Substituting $D=\frac{(N-1)k\theta+N(k-1)}{N-k}$ into the above equation, we obtain the solution:
$\theta=\frac{(N-k)(2N-9n+6)-3nN(k-2)}{(N-3)(N-k)+3kn(N-2)}.$ (3.6)
Then
$F(k,\theta,\beta)=$ $16\left\{\sum\limits_{i=1}^{N}{\beta_{i}^{3}}-\frac{2N-1}{N-1}(1+D)\bar{\lambda}\sum\limits_{i=1}^{N}{\beta_{i}^{2}}\right\}+G(\theta,k){{{\bar{\lambda}}}^{3}}$
$=$ $\frac{16(2N-1)^{3}(1+D)^{3}}{(N-1)^{3}}\bar{\lambda}^{3}\left\{\sum\limits_{i=1}^{N}{\left(\frac{N-1}{(2N-1)(1+D)\bar{\lambda}}{\beta}_{i}\right)^{3}}-\sum\limits_{i=1}^{N}{\left(\frac{N-1}{(2N-1)(1+D)\bar{\lambda}}{\beta}_{i}\right)^{2}}\right\}$
$+G(\theta,k){\bar{\lambda}}^{3}.$
Since
$\sum\limits_{i}{\frac{N-1}{(2N-1)(1+D)\bar{\lambda}}{{\beta}_{i}}}=\frac{N-1}{(2N-1)(1+D)\bar{\lambda}}\sum\limits_{i}{{{\beta}_{i}}}=\frac{N(N-1)}{2N-1},$
by Lemma 3.3 we know the minimal point of $F(k,\theta,\beta)$ is
$(\beta_{1},\beta_{2},\ldots,\beta_{N})=\left((1+D)\bar{\lambda},(1+D)\bar{\lambda},\ldots,(1+D)\bar{\lambda}\right)$
and
$(\beta_{1},\beta_{2},\ldots,\beta_{N})=\left(0,\frac{N(1+D)}{N-1}\bar{\lambda},\ldots,\frac{N(1+D)}{N-1}\bar{\lambda}\right).$
i.e.
$(\lambda_{1},\lambda_{2},\ldots,\lambda_{N})=(\bar{\lambda},\bar{\lambda},\ldots\bar{\lambda})$ (3.7)
and
$(\lambda_{1},\lambda_{2},\ldots,\lambda_{N})=(-D\bar{\lambda},\frac{N+D}{N-1}\bar{\lambda},\ldots,\frac{N+D}{N-1}\bar{\lambda}).$ (3.8)
Hence, the minimal value of $F$ is $F(k,\theta,\bar{\lambda})=0.$
By Lemma 3.1, we have
$\frac{1}{2}\Delta{{\left|R\right|}^{2}}={{\left|\nabla R\right|}^{2}}+\left\langle\Delta R,R\right\rangle\geq 0,$
thus
$\nabla R=0,$
and
$\left\langle\Delta R,R\right\rangle=0.$
Therefore, $M$ is a locally symmetric space. Since when $n\geq 8,$ $\theta(n,k)$ defined by (3.6) satisfies $\theta<\frac{2(n-1)}{n+2}$ for $1\leq k\leq\left[\frac{n+2}{4}\right]$. It follows from Theorem 1.8 and Proposition 2.9 of [Li5] that $M$ is either flat (when $\bar{\lambda}=0$) or a rational homology sphere (when $\bar{\lambda}>0$).
The complete classification of compact symmetric spaces that are rational homology spheres is due to Wolf [Wo, Theorem 1]. Other than spheres, the only simply-connected example is the space SU(3)/SO(3).{SU(3)}/{SO(3)}. (3.7) and (3.8) implies that the eigenvalues of the curvature operator are either
$(\bar{\lambda},\bar{\lambda},\ldots\bar{\lambda})$
or
$(-D\bar{\lambda},\frac{N+D}{N-1}\bar{\lambda},\ldots,\frac{N+D}{N-1}\bar{\lambda}),$
where $D=\frac{(N-1)k\theta+N(k-1)}{N-k}$. In conjunction with the result of [NP, Example 4.5], it follows that the manifold must be a spherical space form, which completes the proof of Theorem A
Remark 3.6.
When $k\leq[\frac{n+2}{4}]$, the expression of $\theta$ in (3.6) must be non-negative, that is
$k\leq\frac{2N(N+3)-3nN}{3n(N-3)+2N+6}\quad\text{when}\quad k\leq[\frac{n+2}{4}].$
When $n\geq 14$, we have
$\frac{2N(N+3)-3nN}{3n(N-3)+2N+6}>\frac{n+2}{4}.$
When $6\leq n\leq 13$,
$\frac{n+2}{4}]<\frac{2N(N+3)-3nN}{3n(N-3)+2N+6}<\frac{n+2}{4},$
except $n=6,7,10$. When $n=6,7$, $k$ can choose $1$ and $n=10$, $k$ can choose $1,2.$
Proof of Theorem B.
When $n=4,5$, according to [DF], due to Jack and Parker’s result [JP] we have
$\left\langle\Delta R,R\right\rangle=8\sum\limits_{i}{\lambda_{i}^{3}}+\frac{8(n-4)}{3}\bar{\lambda}\sum\limits_{i}{\lambda_{i}^{2}}-\frac{4(n+2)(n-1)^{2}}{3}{\bar{\lambda}^{3}}.$
[…]
$\left(-\frac{17}{7}\bar{\lambda},\frac{10}{7}\bar{\lambda},\dots,\frac{10}{7}\bar{\lambda}\right)\quad\text{if}\quad n=4,$
and
$\left(-\frac{47}{18}\bar{\lambda},\frac{23}{18}\bar{\lambda},\dots,\frac{23}{18}\bar{\lambda}\right)\quad\text{if}\quad n=5.$
It follows from [NP, Theorem B] that $M$ is either flat (when $\bar{\lambda}=0$) or a rational homology sphere (when $\bar{\lambda}>0$). The remainder of the argument is analogous to that in Theorem A, which completes the proof of Theorem B. ∎
Proof of Theorem C.
The argument follows the same pattern as in Theorem  A, with (3.6) applied for the cases where $n=6,7$ with $k=1$, and $n=10$ with $k=1,2$. This completes the proof of Theorem C. ∎

### v2

@@
 Abstract.
-Using Bochner techniques, we prove that a compact Einstein manifold of dimension $n\geq 4$ has constant curvature provided that the curvature operator of the second kind satisfies a cone condition that is strictly weaker than nonnegativity.
-Furthermore, employing a result of Li [Li5], we establish that any closed Einstein manifold of dimension $n\geq 4$ satisfying
+Using Bochner techniques, we prove that an Einstein manifold of dimension $n\geq 4$ has constant curvature provided that the curvature operator of the second kind satisfies a cone condition that is strictly weaker than nonnegativity.
+Furthermore, employing a result of Li [Li25], we establish that any closed Einstein manifold of dimension $n\geq 4$ satisfying
 $k^{-1}({\lambda}_{1}+\cdots+{\lambda}_{k})\geq-\theta(n,k)\bar{\lambda},\quad\text{for some}\quad k\leq[\frac{n+2}{4}]$
@@
 Here, ${\lambda}_{1}\leq{\lambda}_{2}\leq\cdots\leq{\lambda}_{\frac{(n-1)(n+2)}{2}}$ are the eigenvalues of $\mathring{R}\,$, $\bar{\lambda}$ is their average, and $\theta(n,k)$ is a positive constant.
-This result generalizes the work of Dai-Fu [DF] and Chen-Wang [CW1, CW].
+This result generalizes the work of Dai-Fu [DF24] and Chen-Wang [CW25, CW24].
+We also classify four-dimensional Einstein manifolds satisfying a cone condition.
 Key words and phrases:
@@
 1. introduction
-The study of the Riemannian curvature operator of the second kind originated with Nishikawa’s conjecture [Ni], which states that a closed simply connected Riemannian manifold with a nonnegative curvature operator of the second kind is diffeomorphic to a locally symmetric space, and if the operator is positive, the manifold is diffeomorphic to a spherical space form.
-This conjecture was confirmed by Cao–Gursky–Tran [CGT] under the assumption of two-positivity, and later strengthened by Li [Li4] and Nienhaus–Petersen–Wink [NP] under the weaker condition of three-nonnegativity.
-Further classification results under various curvature conditions of the second kind can be found in [DFD, Li3, Li7, Li6, Li4, Li2, Li5, DF, NP].
+The study of the Riemannian curvature operator of the second kind originated with Nishikawa’s conjecture [Nis86], which states that a closed simply connected Riemannian manifold with a nonnegative curvature operator of the second kind is diffeomorphic to a locally symmetric space, and if the operator is positive, the manifold is diffeomorphic to a spherical space form.
+This conjecture was confirmed by Cao–Gursky–Tran [CGT23] under the assumption of two-positivity, and later strengthened by Li [Li24a] and Nienhaus–Petersen–Wink [NPW23] under the weaker condition of three-nonnegativity.
+Further classification results under various curvature conditions of the second kind can be found in [DFY24, Li22, Li23a, Li23b, Li24a, Li24b, Li25, DF24, NPW23].
 […]
-Recently, Li [Li5] studied cone conditions of the curvature operator of the second kind, which satisfies
+Recently, Li [Li25] studied cone conditions of the curvature operator of the second kind, which satisfies
 ${\alpha}^{-1}({\lambda}_{1}+\cdots+{\lambda}_{\alpha})\geq-\theta\bar{\lambda},$
-for its eigenvalue of curvature operator of the second kind.
+for the first $\alpha$ smallest eigenvalues of the curvature operator of the second kind.
 This condition is denote by $C(\alpha,\theta)$ and $\mathring{C}\,(\alpha,\theta)$ if the inequality is strict.
 Li proved that if $\mathring{R}\,\in C(\frac{n+2}{2},\theta)$ for some $1<\theta<\frac{2(n-1)}{n+2},$ then $M$ is either flat or a rational homology sphere.
-Building on the work of Li [Li5] and Dai–Fu [DF], Cheng and Wang [CW] relaxed the curvature assumption to a cone condition.
-They introduced a positive constant $\theta(n)$ and proved that if an $n$-dimensional closed Einstein manifold ($n=4,5$, or $n\geq 8$) satisfies $\lambda_{1}+\lambda_{2}\geq-2\theta(n)\bar{\lambda}$, then it is either flat or a spherical space form.
+Building on the work of Li [Li25] and Dai–Fu [DF24], Cheng and Wang [CW24] relaxed the curvature assumption to a cone condition.
+They introduced a positive constant $\theta(n)$ and proved that if an $n(\geq 8)$-dimensional closed Einstein manifold satisfies $\lambda_{1}+\lambda_{2}\geq-2\theta(n)\bar{\lambda}$, then it is either flat or a spherical space form.
 This condition is strictly weaker than $2$-nonnegativity, which corresponds to $\theta=0$.
 In this paper, we investigate Einstein manifolds of dimension $n\geq 4$ whose curvature operator of the second kind satisfies a cone condition that is strictly weaker than nonnegativity.
-Let $\{\lambda_{i}\}_{i=1}^{N}$ be the eigenvalues of $\mathring{R}\,$ and $\bar{\lambda}=\sum\nolimits_{i=1}^{N}{{\lambda}_{i}}/{N}\;$ their average.
-By applying results of Dai–Fu [DF] and Li [Li5], we obtain the following theorems.
+Let $\{\lambda_{i}\}_{i=1}^{N}$ be the eigenvalues of $\mathring{R}\,$ with arranged in non-decreasing order, and $\bar{\lambda}=\sum\nolimits_{i=1}^{N}{{\lambda}_{i}}/{N}\;$ their average.
+By applying results of Dai–Fu [DF24] and Li [Li25], we obtain the following theorems.
 Theorem A.
-Let $(M^{n},g)$ be an $n$-dimensional compact Einstein manifold.
-For $n\geq 8,n\neq 10$ and $k\leq[\frac{n+2}{4}]$, there exists a constant $\theta(n,k)>0$ such that if the curvature operator of the second kind satisfies
+Let $(M^{n},g)$ be an $n$-dimensional Einstein manifold.
+For $n\geq 4,n\neq 6,7,10$ and $k\leq[\frac{n+2}{4}]$, there exists a constant $\theta(n,k)>0$ such that if the curvature operator of the second kind satisfies
 $k^{-1}({\lambda}_{1}+\cdots+{\lambda}_{k})\geq-\theta(n,k)\bar{\lambda},$
@@
 then $M$ is either flat or a spherical space form.
-When $n=4,5$ we have the following theorem.
+Theorem  A admits an enhancement for dimensions $n=4$ and $5$, as stated below.
 Theorem B.
-Let $(M,g)$ be an $n=4$ or $5$-dimensional compact Einstein manifold, and $\mathring{R}\,$ be the curvature operator of the second kind. If
-(i) $n=4$ and $\mathring{R}\,$ satisfies ${\lambda}_{1}\geq-\frac{17}{7}\bar{\lambda}$ or ${\lambda}_{1}+{\lambda}_{2}\geq-\bar{\lambda}$, or
-(ii) $n=5$ and $\mathring{R}\,$ satisfies ${\lambda}_{1}\geq-\frac{47}{18}\bar{\lambda}$ or ${\lambda}_{1}+{\lambda}_{2}\geq-\frac{4}{3}\bar{\lambda}$ or ${\lambda}_{1}+{\lambda}_{2}+{\lambda}_{3}\geq-\frac{1}{18}\bar{\lambda}$,
+Let $(M,g)$ be an $n=4,5$-dimensional Einstein manifold, and $\mathring{R}$ be the curvature operator of the second kind. If
+(i) $n=4$ and $\mathring{R}\,$ satisfies ${\lambda}_{1}\geq-\frac{1}{7}\bar{\lambda}$,
+(ii) $n=5$ and $\mathring{R}\,$ satisfies ${\lambda}_{1}\geq-\frac{4}{9}\bar{\lambda}$,
 then $M$ is either flat or a spherical space form.
 Remark 1.1.
-For dimensions $n=6,7$, we have $\frac{n+2}{4}]=2$, yet our method does not yield a corresponding cone condition for $k=2$.
+For the case $n=4$, Li established a more general cone condition covering all $1\leq k<9$; see [Li25, Throrem 1.5].
+Although this condition is stronger than that of Li, we handle it using the Bochner technique.
+For completeness, we include the details here.
+As will be seen from the subsequent Theorem D and Corollary 4.1, our approach achieves the optimal result.
+Remark 1.2.
+For dimensions $n=6,7$, we have $\frac{n+2}{4}]=2$, but our method does not yield a corresponding cone condition for $k=2$.
 Similarly, for $n=10$, the method does not provide a cone condition for $k=3$.
-Nevertheless, the following theorem remains valid for the cases $n=6,7$ with $k=1$, and for $n=10$ with $k=1,2$.
+Nevertheless, the following theorem remains valid for $n=6,7$ with $k=1$, and for $n=10$ with $k=1,2$.
 Theorem C.
-Let $(M,g)$ be an $n=6,7$ or $10$-dimensional compact Einstein manifold, and $\mathring{R}$ be the curvature operator of the second kind. If
+Let $(M,g)$ be an $n=6,7$ or $10$-dimensional Einstein manifold, and $\mathring{R}$ be the curvature operator of the second kind. If
 (i) $n=6$ and $\mathring{R}\,$ satisfies ${\lambda}_{1}\geq-\frac{208}{647}\bar{\lambda}$,
@@
 Remark 1.3.
-When $\theta=0$, Theorem A,B,C are the result of Dai-Fu [DF]. When $k=1,2$, Theorem A,B,C are the result of Chen-Wang [CW1, CW].
-2. notation and lemmas
-Here, we introduce our notation and summarize key facts regarding the curvature operator of the second kind. Further details can be found in [Li4, Li5, NP, DF, CW].
+When $\theta=0$, Theorem A,B,C are the result of Dai-Fu [DF24].
+When $k=1,2$, Theorem A,B,C are the result of Chen-Wang [CW25, CW24].
+We note that a sign error in [DF24, (4.9)] led to an incorrect Bochner formula for dimensions $4$ and $5$.
+This has now been corrected in (3).
+Hence when $n=4,5$, we can only choose $k=1$.
+Based on Lemma $3.3$, we present a rigidity theorem for Einstein $4$-manifolds in the case $k=3$, which improves [Li25, Theorem 1.8] for $n=4$ and establishes its optimality in this case.
+Theorem D.
+Let $(M,g)$ be an Einstein four-manifold whose the curvature operator of the second kind satisfies
+(1.1) $\lambda_{1}+\lambda_{2}+\lambda_{3}\geq-3\bar{\lambda},$
+then one of the following must be true:
+(1) $M$ is flat.
+(2) up to rescaling, $M$ is isometric to a quotient of the round sphere $(\mathbb{S}^{4},g_{0})$.
+(3) up to rescaling, $M$ is isometric to a $\mathbb{CP}^{2}$ with the Fubini–Study metric.
+(4) up to rescaling, $M$ is isometric to a quotient of $\mathbb{S}^{2}\times\mathbb{S}^{2}$ with the product metric.
+2. preliminaries
+Here, we introduce our notation and summarize key facts regarding the curvature operator of the second kind. Further details can be found in [Li24a, Li25, NPW23, DF24, CW24].
 Let $(V,g)$ be an $n$-dimensional Euclidean vector space and $\left\{e_{i}\right\}_{i=1}^{n}$ be an orthonormal basis for $V$. We recall that the second-order tensor space can be decomposed by
@@
 $\bar{R}({{e}_{i}}\odot{{e}_{j}})=\sum\limits_{k,l}{{{R}_{iklj}}{{e}_{k}}\odot{{e}_{l}}},$
-$\bar{R}$ induces a symmetric form $\mathring{R}\,$ defined by
+where $e_{i}\odot e_{j}=e_{i}\otimes e_{i}+e_{j}\otimes e_{i}.$ $\bar{R}$ induces a symmetric form $\mathring{R}\,$ defined by
 $\mathring{R}\,:S_{0}^{2}(V)\times S_{0}^{2}(V)\to R$
@@
 which is called curvature operator of the second kind.
-It was introduced by Nishikawa [Ni] and as pointed out in [NP], the curvature operator of the second kind $\mathring{R}\,$ can also be interpreted as $\bar{R}$ and its image restrict on $S_{0}^{2}(V)$.
+It was introduced by Nishikawa [Nis86] and as pointed out in [NPW23], the curvature operator of the second kind $\mathring{R}\,$ can also be interpreted as $\bar{R}$ and its image restrict on $S_{0}^{2}(V)$.
 Let $(M,g)$ be an $n$-dimensional Riemannian manifold.
@@
 $\mathring{R}\,$ is denoted by curvature operator of second kind of $M$.
-We introduce the following definition which comes from [NP]
+We introduce the following definition which comes from [NPW23]
 Definition 2.1.
@@
 First we have
-$\sum\limits_{i=1}^{N}{{\lambda}_{i}\left|S^{i}W\right|^{2}}$ $\geq$ $\sum\limits_{i=1}^{k}{{\lambda}_{i}\left|S^{i}W\right|^{2}}+\sum\limits_{i=k+1}^{N}{{\lambda}_{k+1}\left|S^{i}W\right|^{2}}$ (2.1)
+(2.1) $\sum\limits_{i=1}^{N}{{\lambda}_{i}\left|S^{i}W\right|^{2}}$ $\geq$ $\sum\limits_{i=1}^{k}{{\lambda}_{i}\left|S^{i}W\right|^{2}}+\sum\limits_{i=k+1}^{N}{{\lambda}_{k+1}\left|S^{i}W\right|^{2}}$
 $=$ ${{\lambda}_{k+1}}\sum\limits_{i=1}^{N}{\left|S^{i}W\right|^{2}}+\sum\limits_{i=1}^{k}{({\lambda}_{i}-{\lambda}_{k+1})\left|S^{i}W\right|^{2}}$
@@
 If $k^{-1}({\lambda}_{1}+\cdots+{\lambda}_{k})\geq-\theta\bar{\lambda}$, then
-${{\lambda}_{k+1}}\geq\frac{{\lambda}_{1}+\cdots+{\lambda}_{k}}{k}\geq-\theta\bar{\lambda}.$ (2.2)
+(2.2) ${{\lambda}_{k+1}}\geq\frac{{\lambda}_{1}+\cdots+{\lambda}_{k}}{k}\geq-\theta\bar{\lambda}.$
 Applying inequality (2.2) to the right-hand side of (2.1), we get for $0<k\leq[\frac{S}{M}],$
@@
 $=$ $-S\theta\bar{\lambda}.$
-3. proof of main theorem
-Let $(M^{n},g)$ be an Einstein manifold of dimension $n$.
+3. $n$-dimensional Einstein manifolds
+Let $(M^{n},g)$ be an $n$-dimensional Einstein manifold.
 Denote by $\{\lambda_{i}\}_{i=1}^{N}$ the eigenvalues of the operator $\mathring{R}\,$ and let $\bar{\lambda}=\sum\nolimits_{i=1}^{N}{{\lambda}_{i}}/{N}\;$ be their average.
-From [Li4], it follows that
+From [Li24a], it follows that
 $tr(\mathring{R})=\frac{n+2}{2n}s,$
 where $s$ is the scalar curvature of $M$. Consequently $\bar{\lambda}=\frac{s}{n(n-1)}$.
-Based on the calculations presented in [DF], it has been calculated that
-$S=\sum\limits_{i=1}^{N}{{{\left|{{S}^{i}}W\right|}^{2}}}=\frac{2({{n}^{2}}+n-8)}{n}{{\left|W\right|}^{2}}=\frac{4N-12}{n}{{\left|W\right|}^{2}},$ (3.1)
+Based on the calculations presented in [DF24], it has been calculated that
+(3.1) $S=\sum\limits_{i=1}^{N}{{{\left|{{S}^{i}}W\right|}^{2}}}=\frac{2({{n}^{2}}+n-8)}{n}{{\left|W\right|}^{2}}=\frac{4N-12}{n}{{\left|W\right|}^{2}},$
 and
@@
 Moreover, on Einstein manifold, the square norm of the Weyl curvature can be expressed as
-${{\left|W\right|}^{2}}=\frac{4}{3}\sum\limits_{i=1}^{N}{\lambda_{i}^{2}}-\frac{4N}{3}{{\bar{\lambda}}^{2}}.$ (3.2)
+(3.2) ${{\left|W\right|}^{2}}=\frac{4}{3}\sum\limits_{i=1}^{N}{\lambda_{i}^{2}}-\frac{4N}{3}{{\bar{\lambda}}^{2}}.$
 With these preparations we obtain the following estimate for the Bochner formula.
@@
 If $k^{-1}({\lambda}_{1}+\cdots+{\lambda}_{k})\geq-\theta\bar{\lambda}$ for $k\leq[\frac{n+2}{4}],\theta\geq 0$, then the curvature operator $R$ of $M$ satisfies
-$3\left\langle\Delta R,R\right\rangle$ $\geq$ $16\sum\limits_{i=1}^{N}{\lambda_{i}^{3}}+\frac{16}{3n}[2N-12n+6-(N-3)\theta]\bar{\lambda}\sum\limits_{i=1}^{N}{\lambda_{i}^{2}}$ (3.3)
+(3.3) $3\left\langle\Delta R,R\right\rangle$ $\geq$ $16\sum\limits_{i=1}^{N}{\lambda_{i}^{3}}+\frac{16}{3n}[2N-12n+6-(N-3)\theta]\bar{\lambda}\sum\limits_{i=1}^{N}{\lambda_{i}^{2}}$
 $-\frac{16N}{3n}[2N-9n+6-(N-3)\theta]\bar{\lambda}^{3}.$
 Proof.
-The following Bochner formula comes from Proposition 3.4 in [DF].
+The following Bochner formula comes from Proposition 3.4 in [DF24].
 $3\left\langle\Delta R,R\right\rangle=$ $\sum\limits_{i}{{\lambda}_{i}\left|S^{i}W\right|^{2}}+8\left(\frac{-n^{3}+6n^{2}+12n-8}{3n^{4}(n-1)^{2}}\right)s^{3}$
@@
 Remark 3.2.
-When $k=2$, this is Lemma 3.3 of [CW].
+When $k=2$, this is Lemma 3.3 of [CW24].
 To establish the main theorem, we present the following algebraic results, which play a crucial role in the proof.
 Lemma 3.3.
-Let $0\leq{\lambda}_{1}\leq{\lambda}_{2}\leq\cdots\leq{\lambda}_{N}$ and $\sum\limits_{i=1}^{N}{{\lambda}_{i}}=C>0$. If $C=\frac{N(N-1)}{2N-1},$ then the function
+Let $0\leq{\lambda}_{1}\leq{\lambda}_{2}\leq\cdots\leq{\lambda}_{N}$ and $\sum\limits_{i=1}^{N}{{\lambda}_{i}}=C>0$, where $N\geq 3$. If $C=\frac{N(N-1)}{2N-1},$ then the function
 $F(\lambda_{1},\lambda_{2},\ldots,\lambda_{N})=\sum\limits_{i=1}^{N}{\lambda_{i}^{3}}-\sum\limits_{i=1}^{N}{\lambda_{i}^{2}}$
@@
 Case 1. Consider ${\lambda}_{1}>0$. Under the constraints $\sum\limits_{i}{{{\lambda}_{i}}}=C>0$, Lagrange multipliers yields a constant $\mu\in\mathbb{R}$ such that
-$3\lambda_{i}^{2}-2{\lambda}_{i}+\mu=0$ (3.4)
+(3.4) $3\lambda_{i}^{2}-2{\lambda}_{i}+\mu=0$
 for $i=1,\ldots,N.$ The possible solution of this equation is
@@
 Because $C=\frac{N(N-1)}{2N-1}$ implies 3CN>03C-N>0, we must have N2k>0N-2k>0. Consequently,
-$\mu=\frac{1}{3}[1-{\left(\frac{3C-N}{N-2k}\right)^{2}}].$ (3.5)
+(3.5) $\mu=\frac{1}{3}[1-{\left(\frac{3C-N}{N-2k}\right)^{2}}].$
 Substituting $Q_{k}$ into $F$ and applying the relation (3.4) to obtain
@@
 So in $\partial\Omega,$
-$\underset{\partial\Omega}{\mathop{\min}}\,F=\underset{1\leq k<N}{\mathop{\min}}\,\frac{C^{3}}{(N-k)^{2}}-\frac{C^{2}}{N-k}.$
+$\underset{\partial\Omega}{\mathop{\min}}\,F=\underset{1\leq k<N}{\mathop{\min}}\,\{\frac{C^{3}}{(N-k)^{2}}-\frac{C^{2}}{N-k}\}.$
 Since we assume $C=\frac{N(N-1)}{2N-1},$ we obtain
@@
 equality holds if and only if $k=1$. Hence,
-$\underset{\partial\Omega}{\mathop{\min}}\,F=\underset{1\leq k<N}{\mathop{\min}}\,\frac{C^{3}}{(N-k)^{2}}-\frac{C^{2}}{N-k}=\frac{C^{3}}{N^{2}}-\frac{C^{2}}{N},$
+$\underset{\partial\Omega}{\mathop{\min}}\,F=\underset{1\leq k<N}{\mathop{\min}}\,\{\frac{C^{3}}{(N-k)^{2}}-\frac{C^{2}}{N-k}\}=\frac{C^{3}}{N^{2}}-\frac{C^{2}}{N},$
 and the minimal point is $\left(0,\frac{C}{N-1},\ldots,\frac{C}{N-1}\right)$. As $X$ possesses at least one zero, it follows that the minimizer must be unique.
@@
 Proof of Theorem A.
-If $\bar{\lambda}=0$, then $M$ is flat. Now we consider $\bar{\lambda}>0$.
+According to [Li25, Proposition 2.8], we have $\bar{\lambda}\geq 0$ for $\theta(n,k)>0$.
+If $\bar{\lambda}=0$, then $M$ is flat.
+Now we consider the case $\bar{\lambda}>0$, i.e., $s>0$.
+Hence $M$ is compact by Myers Theorem.
 Writing the right term of (3.3) as a function $F(k,\theta,\lambda)$,
@@
 Substituting $D=\frac{(N-1)k\theta+N(k-1)}{N-k}$ into the above equation, we obtain the solution:
-$\theta=\frac{(N-k)(2N-9n+6)-3nN(k-2)}{(N-3)(N-k)+3kn(N-2)}.$ (3.6)
+(3.6) $\theta=\frac{(N-k)(2N-9n+6)-3nN(k-2)}{(N-3)(N-k)+3kn(N-2)}.$
 Then
@@
 i.e.
-$(\lambda_{1},\lambda_{2},\ldots,\lambda_{N})=(\bar{\lambda},\bar{\lambda},\ldots\bar{\lambda})$ (3.7)
+(3.7) $(\lambda_{1},\lambda_{2},\ldots,\lambda_{N})=(\bar{\lambda},\bar{\lambda},\ldots\bar{\lambda})$
 and
-$(\lambda_{1},\lambda_{2},\ldots,\lambda_{N})=(-D\bar{\lambda},\frac{N+D}{N-1}\bar{\lambda},\ldots,\frac{N+D}{N-1}\bar{\lambda}).$ (3.8)
+(3.8) $(\lambda_{1},\lambda_{2},\ldots,\lambda_{N})=(-D\bar{\lambda},\frac{N+D}{N-1}\bar{\lambda},\ldots,\frac{N+D}{N-1}\bar{\lambda}).$
 Hence, the minimal value of $F$ is $F(k,\theta,\bar{\lambda})=0.$
@@
 Since when $n\geq 8,$ $\theta(n,k)$ defined by (3.6) satisfies $\theta<\frac{2(n-1)}{n+2}$ for $1\leq k\leq\left[\frac{n+2}{4}\right]$.
-It follows from Theorem 1.8 and Proposition 2.9 of [Li5] that $M$ is either flat (when $\bar{\lambda}=0$) or a rational homology sphere (when $\bar{\lambda}>0$).
-The complete classification of compact symmetric spaces that are rational homology spheres is due to Wolf [Wo, Theorem 1].
+It follows from Theorem 1.8 and Proposition 2.9 of [Li25] that $M$ is either flat (when $\bar{\lambda}=0$) or a rational homology sphere (when $\bar{\lambda}>0$).
+The complete classification of compact symmetric spaces that are rational homology spheres is due to Wolf [Wol69, Theorem 1].
 Other than spheres, the only simply-connected example is the space SU(3)/SO(3).{SU(3)}/{SO(3)}.
@@
 $(-D\bar{\lambda},\frac{N+D}{N-1}\bar{\lambda},\ldots,\frac{N+D}{N-1}\bar{\lambda}),$
-where $D=\frac{(N-1)k\theta+N(k-1)}{N-k}$. In conjunction with the result of [NP, Example 4.5], it follows that the manifold must be a spherical space form, which completes the proof of Theorem A
+where $D=\frac{(N-1)k\theta+N(k-1)}{N-k}$. In conjunction with the result of [NPW23, Example 4.5], it follows that the manifold must be a spherical space form, which completes the proof of Theorem A
 Remark 3.6.
@@
 $\frac{2N(N+3)-3nN}{3n(N-3)+2N+6}>\frac{n+2}{4}.$
-When $6\leq n\leq 13$,
+When $4\leq n\leq 13$,
 $\frac{n+2}{4}]<\frac{2N(N+3)-3nN}{3n(N-3)+2N+6}<\frac{n+2}{4},$
@@
 Proof of Theorem B.
-When $n=4,5$, according to [DF], due to Jack and Parker’s result [JP] we have
-$\left\langle\Delta R,R\right\rangle=8\sum\limits_{i}{\lambda_{i}^{3}}+\frac{8(n-4)}{3}\bar{\lambda}\sum\limits_{i}{\lambda_{i}^{2}}-\frac{4(n+2)(n-1)^{2}}{3}{\bar{\lambda}^{3}}.$
-[…]
-$\left(-\frac{17}{7}\bar{\lambda},\frac{10}{7}\bar{\lambda},\dots,\frac{10}{7}\bar{\lambda}\right)\quad\text{if}\quad n=4,$
-and
-$\left(-\frac{47}{18}\bar{\lambda},\frac{23}{18}\bar{\lambda},\dots,\frac{23}{18}\bar{\lambda}\right)\quad\text{if}\quad n=5.$
-It follows from [NP, Theorem B] that $M$ is either flat (when $\bar{\lambda}=0$) or a rational homology sphere (when $\bar{\lambda}>0$).
+The corrected version of [DF24, (4.9)] is as follows.
+${R_{tspq}}{R_{tjpl}}{R_{sjql}}=$ ${W_{tspq}}{W_{tjpl}}{W_{sjql}}-\frac{3s}{2n(n-1)}\left|W\right|^{2}+\frac{(n-2)s^{3}}{n^{2}(n-1)^{2}}$
+$=$ ${W_{tspq}}{W_{tjpl}}{W_{sjql}}-\frac{3s}{2n(n-1)}\left|R\right|^{2}+\frac{(n+1)s^{3}}{n^{2}(n-1)^{2}}.$
+Here the sign for $\frac{{3s}}{{2n(n-1)}}{\left|W\right|^{2}}$ is given as ‘++’ in [DF24, (4.9)], but it should be corrected to ‘-’.
+Following the approach of [DF24], we present the corrected formula as follows:
+$\left\langle\Delta R,R\right\rangle=8\sum\limits_{i}{\lambda_{i}^{3}}+\frac{8(n-10)}{3n(n-1)}s\sum\limits_{i}{\lambda_{i}^{2}}-\frac{4(n-7)(n+2)}{3n^{3}(n-1)^{2}}s^{3}$
+(3.9) $=$ $8\sum\limits_{i}{\lambda_{i}^{3}}+\frac{8(n-10)}{3}\bar{\lambda}\sum\limits_{i}{\lambda_{i}^{2}}-\frac{4(n+2)(n-7)(n-1)}{3}\bar{\lambda}^{3}.$
+We use the above Bochner formula, and the same methods as the proof of Theorem A, then take $n=4,5$ and $k=1$.
+We get $n=4$, $\theta=\frac{1}{7}$ and $n=5,\theta=\frac{4}{9}$.
 The remainder of the argument is analogous to that in Theorem A, which completes the proof of Theorem B.
@@
 The argument follows the same pattern as in Theorem  A, with (3.6) applied for the cases where $n=6,7$ with $k=1$, and $n=10$ with $k=1,2$. This completes the proof of Theorem C. ∎
+4. $4$-dimensional Einstein manifolds
+In this section, we present a geometric application of Lemma 3.3 in $4$-dimensional Einstein manifolds.
+Proof of Theorem D.
+The Bochner formula (3) on $4$-dimensional Einstein manifolds can be written as
+(4.1) $\left\langle\Delta R,R\right\rangle=8\sum\limits_{\alpha}{\lambda_{\alpha}^{3}}-\frac{4}{3}s\sum\limits_{\alpha}{\lambda_{\alpha}^{2}}+\frac{1}{24}s^{3}.$
+Denote the eigenvalues of $W^{+}$ by $a_{1}\leq a_{2}\leq a_{3}$, and those of $W^{-}$ by $b_{1}\leq b_{2}\leq b_{3}$, then
+(4.2) $a_{1}+a_{2}+a_{3}=0,b_{1}+b_{2}+b_{3}=0.$
+According to [CGT23], on Einstein 4-manifold, the eigenvalues of $\mathring{R}$ are given by
+(4.3) ${\lambda}_{ij}=\frac{s}{12}-a_{i}-b_{j},\quad i,j=1,2,3.$
+If $\mathring{R}$ satisfies (1.1), then
+$a_{3}=\frac{s}{12}-\frac{\lambda_{31}+\lambda_{32}+\lambda_{33}}{3}\leq\frac{s}{12}+\bar{\lambda}=\frac{s}{6},$
+and
+$b_{3}=\frac{s}{12}-\frac{\lambda_{13}+\lambda_{23}+\lambda_{33}}{3}\leq\frac{s}{12}+\bar{\lambda}=\frac{s}{6}.$
+Consequently, $s\geq 0$; moreover, if $s=0$, then $W^{\pm}=0$, i.e., $M$ is flat.
+By substituting (4.3) into (4.1) and applying (4.2), we directly compute to obtain
+(4.4) $\left\langle\Delta R,R\right\rangle=2[s(a_{1}^{2}+a_{2}^{2}+a_{3}^{2})-12(a_{1}^{3}+a_{2}^{3}+a_{3}^{3})]+2[s(b_{1}^{2}+b_{2}^{2}+b_{3}^{2})-12(b_{1}^{3}+b_{2}^{3}+b_{3}^{3})].$
+Let
+$f_{1}(a_{1},a_{2},a_{3})=s(a_{1}^{2}+a_{2}^{2}+a_{3}^{2})-12(a_{1}^{3}+a_{2}^{3}+a_{3}^{3}),f_{2}(b_{1},b_{2},b_{3})=s(b_{1}^{2}+b_{2}^{2}+b_{3}^{2})-12(b_{1}^{3}+b_{2}^{3}+b_{3}^{3}).$
+Now we consider the case $s>0$. The expression for $f_{1}$ can be rewritten as
+$f_{1}=s(a_{1}^{2}+a_{2}^{2}+a_{3}^{2})-12(a_{1}^{3}+a_{2}^{3}+a_{3}^{3})$
+$=s\sum\limits_{i=1}^{3}{(\frac{s}{6}-a_{i}-\frac{s}{6})^{2}}+12\sum\limits_{i=1}^{3}{(\frac{s}{6}-a_{i}-\frac{s}{6})^{3}}$
+$=12\sum\limits_{i=1}^{3}{(\frac{s}{6}-a_{i})^{3}}-5s\sum\limits_{i=1}^{3}{(\frac{s}{6}-a_{i})^{2}}+\frac{s^{3}}{4}$
+$=\frac{125}{144}s^{3}\left[\sum\limits_{i=1}^{3}{[\frac{12}{5s}(\frac{s}{6}-a_{i})]^{3}}-\sum\limits_{i=1}^{3}{[\frac{12}{5s}(\frac{s}{6}-a_{i})]^{2}}\right]+\frac{s^{3}}{4}.$
+Since
+$\frac{12}{5s}(\frac{s}{6}-a_{i})\geq 0\quad\text{and}\quad\sum\limits_{i}{\frac{12}{5s}(\frac{s}{6}-a_{i})}=\frac{6}{5},$
+it follows from Lemma 3.3 that the minimal points of $f_{1}$ are
+$(\frac{12}{5s}(\frac{s}{6}-a_{1}),\frac{12}{5s}(\frac{s}{6}-a_{2}),\frac{12}{5s}(\frac{s}{6}-a_{3}))=(\frac{2}{5},\frac{2}{5},\frac{2}{5})$
+and
+$(\frac{12}{5s}(\frac{s}{6}-a_{1}),\frac{12}{5s}(\frac{s}{6}-a_{2}),\frac{12}{5s}(\frac{s}{6}-a_{3}))=(\frac{3}{5},\frac{3}{5},0),$
+which correspond respectively to
+$(a_{1},a_{2},a_{3})=(0,0,0)\quad\text{and}\quad(a_{1},a_{2},a_{3})=(-\frac{s}{12},-\frac{s}{12},\frac{s}{6}).$
+Hence,
+$f_{1}\geq f_{1}(0,0,0)=f_{1}(-\frac{s}{12},-\frac{s}{12},\frac{s}{6})=0.$
+Similarly, the minimal points of $f_{2}$ are $(b_{1},b_{2},b_{3})=(0,0,0)$ and $(b_{1},b_{2},b_{3})=(-\frac{s}{12},-\frac{s}{12},\frac{s}{6}),$ and
+$f_{2}\geq f_{2}(0,0,0)=f_{2}(-\frac{s}{12},-\frac{s}{12},\frac{s}{6})=0.$
+In summary, we have
+$\frac{1}{2}\Delta{{\left|R\right|}^{2}}={{\left|\nabla R\right|}^{2}}+\left\langle\Delta R,R\right\rangle\geq 0.$
+By the maximum principle, we obtain $\nabla R=0$, and either
+[…]
+[…]
+[…]
+[…]
+Therefore, the curvature operator $\hat{R}$ of the first kind has eigenvalues $(\frac{s}{12},\frac{s}{12},\frac{s}{12},\frac{s}{12},\frac{s}{12},\frac{s}{12})$, $(\frac{s}{12},\frac{s}{12},\frac{s}{12},0,0,\frac{s}{4})$, $(0,0,\frac{s}{4},\frac{s}{12},\frac{s}{12},\frac{s}{12})$, or $(0,0,\frac{s}{4},0,0,\frac{s}{4})$.
+Applying the classification of $4$-dimensional symmetric spaces, we complete the proof of Theorem D.
+∎
+Corollary 4.1.
+Let $(M,g)$ be an Einstein four-manifold with $4\frac{1}{2}$-nonnegative curvature operator of the second kind. Then one of the following must be true:
+(1) $M$ is flat.
+(2) up to rescaling, $M$ is isometric to a quotient of the round sphere $(\mathbb{S}^{4},g_{0})$.
+(3) up to rescaling, $M$ is isometric to $\mathbb{CP}^{2}$ with the Fubini–Study metric.
+Remark 4.2.
+This corollary has already been proven in [CGT23] using a different method.
+Proof.
+Suppose that ${\lambda}_{1}\leq{\lambda}_{2}\leq\cdots\leq{\lambda}_{9}$ are the eigenvalues of $\mathring{R}\,$ and $\bar{\lambda}$ is their average.
+Then $4\frac{1}{2}$-nonnegative curvature operator of the second kind $\mathring{R}\,$ satisfies
+${\lambda}_{1}+{\lambda}_{2}+{\lambda}_{3}\geq-{\lambda}_{4}-\frac{1}{2}{\lambda}_{5}$
+$\geq-\frac{\frac{1}{2}{\lambda}_{5}+{\lambda}_{6}+{\lambda}_{7}+{\lambda}_{8}+{\lambda}_{9}}{3}$
+$\geq-\frac{{\lambda}_{1}+{\lambda}_{2}+{\lambda}_{3}+{\lambda}_{4}+{\lambda}_{5}+{\lambda}_{6}+{\lambda}_{7}+{\lambda}_{8}+{\lambda}_{9}}{3}$
+$=-3\bar{\lambda}.$
+Combining Theorem D, we complete the proof of Corollary 4.1 because $\mathbb{S}^{2}\times\mathbb{S}^{2}$ has $4\frac{1}{2}$-negative curvature operator of the second kind. ∎